OpenStudy (anonymous):

What is the simplified form of the expression? k^3(k7/5)^-5 how do i solve this?

OpenStudy (paxpolaris):

is this right? \[\huge k^3\cdot \left(k^{7 /5}\right)^{-5}\]

OpenStudy (paxpolaris):

\[\Large \left( a^m \right)^n=a^{m n}\] so, we can simplify it to: \[\huge =k^3\cdot \left(k^{\left( \frac7 {\cancel5} \cdot(-\cancel5) \right) }\right) \] \[\huge =k^3\cdot k^{-7} = k^{-4}\] \[\Huge ={1 \over k^4}\]